OpenStudy (anonymous):

A ball is thrown directly upward from ground level with an initial speed of 80 ft/s. How high will it go? When will it return to the ground? explain please.

OpenStudy (anonymous):

Do you know the acceleration of gravity on Earth's surface?

OpenStudy (anonymous):

it doesnt say in the problem.

OpenStudy (anonymous):

you can solve it by conservation of energy... 1/2*v^2 = g*h Although you exclude drag from the air in this equation.

OpenStudy (anonymous):

Well, the problem probably assumes you know the acceleration of gravity, which is -32f/s^s. So plug that into the motion equation v^2 = v0^2 + 2as, and solve for s.

OpenStudy (anonymous):

1/2*80^2 = 32*h h = 100 ft